OpenStudy (anonymous):

solve 4^3p(4^2p-1=100

OpenStudy (anonymous):

the p goes along with the exponents on both sides

OpenStudy (australopithecus):

so, \[4^{3p}(4^{2p} - 1) = 100\]

OpenStudy (anonymous):

also add the -1 to the second part

OpenStudy (anonymous):

i ment along with the exponent 2p

OpenStudy (australopithecus):

ugh equation editor isn't working anymore

OpenStudy (australopithecus):

also ignore the last line I made a mistake and forgot to add a p

OpenStudy (australopithecus):

I dont know if I'm on the right track to solving it but I can't attempt to complete my solution because this site is buggy

OpenStudy (anonymous):

well i am lost, all i know is that the aswer has to be 0.8644

OpenStudy (australopithecus):

Oh I made a stupid mistake

OpenStudy (australopithecus):

\[4^{3p}(4^{2p-1}) = 100\] \[4^{3p+ 2p - 1} = 100\] \[4^{5p - 1} = 100\] \[4^{5p}4^{- 1} = 100\] \[\frac{4^{5p}}{4^{1}} = 100\] \[\frac{4^{5p}}{4} = 100\] \[4^{5p} = 100(4)\] \[\log_4{4^{5p}} = \log_4{100(4)}\] \[5p = \log_4{100(4)}\] \[p = \frac{\log_4{100(4)}}{5}\]

OpenStudy (australopithecus):

there is your solution

OpenStudy (australopithecus):

I wish I was better at algebra so I could solve that other problem but it probably required some crazy technique ha

OpenStudy (anonymous):

thank you so much this problem was making me mad

OpenStudy (australopithecus):

well it isn't that hard of a problem, it is just application of some fundamental rules of logarithms and exponents

OpenStudy (australopithecus):

do you follow how I came to the solution?

OpenStudy (australopithecus):

If you do not understand part of my solution tell me which part and I will show you the rule I applied

OpenStudy (anonymous):

to be honest no

OpenStudy (australopithecus):

do you not understand any of it?

OpenStudy (australopithecus):

I'm about to go to sleep so please be quick in your responses

OpenStudy (anonymous):

its better u go to bed, i will just ask i class tomorrow. thanks again

OpenStudy (australopithecus):

Well I can explain the rules to you if you want

OpenStudy (australopithecus):

unless you take me for a poor teacher ha which I'm fine with